Information on Grand Corporation’s direct materials costs follows.

Quantities of chemical Y purchased and used 19,400 gallons
Actual cost of chemical Y used $ 426,000
Standard price per gallon of chemical Y $ 22.70
Standard quantity of chemical Y allowed 17,800 gallons

Grand Corporation has no materials inventories.

Required:

a. What were Grand Corporation’s direct materials price and efficiency variances?
b. (Appendix) Prepare the journal entries to record the purchase and use of chemical Y using standard costing.
Answer #1
GRAND CORPORATION
a.
Quantity Price Cost
Actual 19400 $          21.96 $                                                         426,000
Standard 17800 $          22.70 $                                                         404,060
Material Price Variance = (Actual Price- Standard Price)* Actual Qty
= (21.96-22.70)*19400
= 14380 Favourable
Material Efficiency Variance = (Actual Qty- Standard Qty)*Standard Price
= (19400-17800)*22.70
= 36320 Unfavourble
Material Cost Variance = Actual Cost-Standard Cost
= 426000-404060
= 21940 Unfavourble
Material Cost Variance = Material Price Variance + Material Efficiency Variance
= 14380 (F) + 36320 (UF)
= 21940 Unfavourble
b.
Journal Entries
Date Particulars Debit Credit
Chemical Y (17800 gallons * $22.70/ per gallon) 404060
Direct Material Cost Variance 21940
To Accounts Payable 426000
Chemical Y (FG) (19400* $22.70/per gallon) 440380
To Direct Material Efficiency Variance 36320
To Chemical Y (17800 gallons * $22.70/ per gallon) 404060